### v2

# Two-stage Design for Failure Probability Estimation with Gaussian Process Surrogates

Abstract
We tackle the problem of quantifying failure probabilities for expensive deterministic computer experiments with stochastic inputs. The computational cost of the computer simulation prohibits direct Monte Carlo (MC) and necessitates a surrogate model, turning the problem into a two-stage enterprise (surrogate training followed by probability estimation). Limited budgets create a design problem: how should expensive evaluations be allocated between and within the training and estimation stages? One may use the entire evaluation budget to sequentially train the surrogate through contour location (CL), with failure probabilities then estimated solely from the surrogate (we call it “surrogate MC”). But extended CL offers diminishing returns, and surrogate MC relies too stringently on surrogate accuracy. Alternatively, a partially trained surrogate may inform importance sampling, but this can provide erroneous results when budgets are limited. Instead we propose a two-stage design: starting with sequential CL, halting CL once learning has plateaued, then greedily allocating the remaining budget to MC samples with high classification entropy. Ultimately, we employ a “hybrid MC” estimator which leverages the trained surrogate in conjunction with the true responses observed in this second stage. Our unique two-stage design strikes an appropriate balance between exploring and exploiting, and outperforms alternatives, including both of the aforementioned approaches, on a variety of benchmark exercises. With these tools, we are able to effectively estimate small failure probabilities with only hundreds of simulator evaluations, showcasing functionality with both shallow and deep Gaussian process surrogates, and deploying our method on a simulation of fluid flow around an airfoil.
Keywords: computer experiment, contour location, deep Gaussian process, emulator, entropy, importance sampling, Monte Carlo
1 Introduction
Quantifying the probability of failure in complex dynamic systems is of great interest but is challenging due to the computational expense of the so-called “computer experiment.” Consider a deterministic black-box computer simulation, $f:\mathcal{X}\rightarrow\mathbb{R}$ with inputs $\mathcal{X}\subset\mathbb{R}^{d}$ governed by known measure $\mathbb{P}_{\mathbf{x}}$ (with density $p(\mathbf{x})$). The response is thresholded such that $f(\mathbf{x})>t$ indicates failure. The research objective is to estimate the failure probability,
[…]
We are particularly motivated by aeronautic applications where system failures may incur immense financial or environmental costs, or even loss of life. In aircraft design and development, computer simulation experiments are a natural, often mandatory, precursor to direct experimentation and prototyping. For example, $f(\cdot)$ may represent the vibrations experienced by an aircraft which become unsafe above certain levels, or the fuel efficiency of an aircraft, which if too low, can be detrimental to the environment, obliterate fuel budgets, or overly restrict travel distances. Inputs $\mathcal{X}$ may encapsulate design parameters and operating conditions (such as wing shapes, flight speeds, and angles of attack). Yet we acknowledge a slew of other relevant applications where quantifying failure probabilities for complex simulations is critical, including structural reliability analysis using finite element simulations (e.g., Mia et al.,, 2017), Boltzmann modeling of chemical processes in fuel filtration (e.g., Belot et al.,, 2021), and thermodynamic flows for nuclear reactors (e.g., Batet et al.,, 2014).
A rudimentary approach to solve Eq. (1) involves brute force Monte Carlo (MC): simply sample an exhaustively large set of inputs, evaluate the computer simulation at each of these configurations, and report the proportion that resulted in failure. But this approach is prohibitively expensive if the computer simulation is costly (incurring long wall clock times), the probability of failure is rare (requiring enormous sample sizes), or both! In this work, we propose a new strategy for quantifying failure probabilities of complex computer simulations when the computation budget restricts evaluations of the simulator to only a few hundred, rendering traditional MC impractical.
In the face of expensive data, statistical models ($\hat{f}$, called “surrogates” or “emulators”) are essential. Surrogates are trained on a limited set of simulator evaluations in order to provide predictions with appropriate uncertainty quantification (UQ) at unobserved inputs. Gaussian processes (GPs) are the canonical choice (Santner et al.,, 2018; Gramacy,, 2020), but deep Gaussian processes (DGPs; Damianou and Lawrence,, 2013) are becoming increasingly popular, particularly for nonstationary computer experiments (Sauer,, 2023). Notably, our contribution is agnostic to surrogate choice as long as effective UQ is provided, and is thus applicable to traditional GPs and Bayesian DGPs (Sauer et al.,, 2023b) and, potentially, Bayesian neural networks (MacKay,, 1995).
Choosing a surrogate model, like a GP or DGP, is only a first step. The utility of a surrogate boils down to two things: how it is trained and how it is used. For our objective (1), surrogate training should prioritize accurate classification of failures and surrogate deployment should focus on effective estimation of $\alpha$. The prevailing question is then: how can a limited budget of expensive simulator evaluations be optimally allocated between and within the training and estimation stages?
We briefly digress to acknowledge some early methods of failure probability estimation, namely the first and second order reliability methods (FORM/SORM; Haldar and Mahadevan,, 1995), which leveraged simplistic linear approximations of the complex model and placed all failure probability estimation in the hands of the “most probable point” (MPP). Although advancements to FORM and SORM have been made with more flexible GP models (e.g., Zhang et al.,, 2015; Su et al.,, 2017), sole reliance on a MPP is too restricting for our target applications.
To solve Eq. (1), surrogate accuracy in classifying failures is of primary importance. Not all training designs are created equally. Space-filling designs (Joseph,, 2016) may miss failure regions altogether, particularly when training budgets are limited. Instead, it is common to train a surrogate on a small initial design then pair it with a relevant acquisition function to select subsequent inputs for evaluation through the simulator. This iterative process of acquisition and surrogate updating can strategically hone in on the contour, all while minimizing evaluations of the expensive simulator. So-called “contour location” (CL) naturally outperforms space-filling counterparts. Acquisition functions seek to balance exploitation and exploration by leveraging both predicted distance from the contour and posterior uncertainty. Popular variations fall under expected improvement frameworks (Bichon et al.,, 2008; Ranjan et al.,, 2008), stepwise uncertainty reduction frameworks (Bect et al.,, 2012; Chevalier et al.,, 2014; Azzimonti et al.,, 2021), or both (Duhamel et al.,, 2023). Others utilize pass/fail classification entropy (Marques et al.,, 2018; Cole et al.,, 2023). Booth et al., (2024b) recently extended CL sequential designs to DGP surrogates. For small failure regions, contour location may be replaced with subset simulation (Au and Beck,, 2001) which initially targets a conservative threshold ($t^{\star}<t$) and iteratively refines it until the true threshold is found ($t^{\star}\rightarrow t$) (Dubourg et al.,, 2011; Huang et al.,, 2016; Hristov et al.,, 2019). Yet these approaches require Markov chain Monte Carlo (MCMC) sampling of cascading input distributions, which we intentionally avoid here.
A well trained surrogate may be leveraged directly within an MC estimator, with surrogate predictions (which are of neglible cost) in place of true simulator evaluations. One could reasonably use the entire budget of simulator evaluations for contour location, then deploy the trained surrogate in such a “surrogate MC” estimator. Although this is common (e.g., Zhu and Du,, 2016; Cheng and Lu,, 2020; Li et al.,, 2021; Lu et al.,, 2023), we find that surrogate learning of failure contours offers quickly diminishing returns. Furthermore, surrogate MC estimators place too much stock in the accuracy of the surrogate and funnel surrogate inaccuracies directly into the failure probability estimate.
Another common approach splits the budget of simulator evaluations into two disjoint stages: some evaluations are used to train the surrogate and some are used for importance sampling (IS; Srinivasan,, 2002; Tokdar and Kass,, 2010). IS hinges on the construction of a bias distribution that intentionally focuses density in the failure region. Inputs are sampled from the bias distribution and are then utilized in a weighted average calculation that corrects for the disparity between the bias density and the original input density. Identifying an effective bias distribution in the face of limited training data is a hefty task (Tabandeh et al.,, 2022). Adaptive methods abound (e.g., Oh and Berger,, 1992; Kurtz and Song,, 2013; Dalbey and Swiler,, 2014; Bugallo et al.,, 2017; Cheng et al.,, 2023). Peherstorfer et al., (2016) propose what they termed “multifidelity importance sampling,” although we prefer the term “surrogate-informed importance sampling” (SIIS) to avoid confusion with multifidelity surrogate modeling (e.g., Park et al.,, 2017). In SIIS, a trained surrogate is used to identify predicted failures, and a bias distribution is fit to these locations (more on this in Section 2.3). SIIS has been deployed with GP surrogates in a variety of subsequent works (e.g., Cole et al.,, 2023; Renganathan et al.,, 2023; Booth et al.,, 2024a; Renganathan,, 2024), and we view it as our primary competitor. Yet SIIS has several crucial drawbacks. We will delay a thorough discussion to Section 2.3, but the highlights are: (i) SIIS relies too heavily on an arbitrary bias distribution, (ii) it largely ignores the carefully trained surrogate, and (iii) it offers no guidance for allocating the budget between surrogate training and IS estimation.
We prefer to avoid a bias distribution altogether, instead focusing on how we could optimally leverage a surrogate and an MC estimator in tandem. The answer lies in a two-stage design. First, we leverage sequential contour location to pin down the failure contour. At the start, it is essential to target inputs which provide the most insight into the entire failure contour (getting the most “bang for our buck” from early observations). As training progresses, successive acquisitions will offer less additional insight, indicated by apparent convergence of the surrogate MC estimator. Although we can not be sure of theoretical convergence with our limited budget, we propose a decision rule to recognize practical convergence (when surrogate learning from CL has saturated). We then spend the remaining simulation budget precisely on the Monte Carlo samples with the highest classification entropy, ultimately training the surrogate on observations from both stages. Our final failure probability estimate is a “hybrid MC” estimator which combines the observations from this second stage with surrogate predictions at unobserved samples.
This manuscript is laid out as follows. Section 2, reviews GP/DGP surrogates, contour location, and importance sampling. Section 3.2 presents our two-stage design and failure probability estimator. Section 4 presents a variety of synthetic benchmarks. Finally, we deploy our method on a computer experiment of an RAE-2822 airfoil in Section 5 and conclude in Section 6.
2 Preliminaries
Denote single inputs as the row vector $\mathbf{x}$ of size $1\times d$ with response $y=f(\mathbf{x})$. Let $X_{n}$ represent the row-combined matrix of $n$-many inputs with corresponding response vector $\mathbf{y}_{n}$. Our objective is to quantify the failure probability from Eq. (1), namely $\mathrm{P}(f(\mathbf{x})>t)$ for $\mathbf{x}\sim p(\mathbf{x})$. Typical Monte Carlo with a budget of size $M$ provides the estimator
[…]
This estimator is ineffective unless a sufficient number of failures are observed, necessitating very large $M$ particularly for small $\alpha$. With an expensive computer model, we simply cannot evaluate $f(\mathbf{x})$ thousands/millions/billions of times. Throughout, let $N$ denote the total budget of expensive simulator evaluations alotted. We will describe our sample efficient method for estimating failure probabilities in Section 3.2, targeting $N$ in the hundreds, after a quick review of preliminaries.
2.1 Gaussian process surrogates
A surrogate model $\hat{f}$, trained on a limited budget of simulations, may provide predictions at unobserved inputs in place of the true $f$. To enable our proposed methodology, the surrogate must also provide posterior predictive distributions. Gaussian processes (GPs) are the predominant choice; they assume a multivariate normal distribution over the response. For any $n$ observed locations, our GP prior assumption is $\mathbf{y}_{n}\sim\mathcal{N}(\mu,\Sigma(X_{n}))$. The prior mean $\mu$ is often simplified to the zero vector after centering (which we will use moving forward), and the prior covariance $\Sigma(X_{n})$ is typically a function of inverse euclidean distances, e.g., $\Sigma(X_{n})^{ij}=\Sigma(\mathbf{x}_{i},\mathbf{x}_{j})=k(||\mathbf{x}_{i}-\mathbf{x}_{j}||^{2})$ with kernel $k(\cdot)$. See Santner et al., (2018); Rasmussen et al., (2006); Gramacy, (2020) for thorough reviews. Conditioned on training data $\left\{X_{n},\mathbf{y}_{n}\right\}$, posterior predictions at singleton input $\mathbf{x}$ follow
$\hat{f}_{n}(\mathbf{x})\sim\mathcal{N}\left(\mu_{n}(\mathbf{x}),\sigma_{n}^{2}(\mathbf{x})\right)\quad\textrm{for}\quad\begin{array}[]{rl}\mu_{n}(\mathbf{x})&=\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\mathbf{y}_{n}\\ \sigma_{n}^{2}(\mathbf{x})&=\Sigma(\mathbf{x},\mathbf{x})-\Sigma(\mathbf{x},X_{n})\Sigma(X_{n})^{-1}\Sigma(\mathbf{x},X_{n})^{\top},\end{array}$ (3)
where $\Sigma(\mathbf{x},X_{n})$ denotes the vector formed from applying the kernel between $\mathbf{x}$ and every row of $X_{n}$. Joint predictions are straightforward but not necessary for our contribution.
A natural alternative to true MC (2) is “surrogate MC” which leverages surrogate predictions in place of the true simulator,
[…]
Here we use the posterior mean of our GP in place of $f$. If the surrogate is not accurate near the contour, it will misclassify inputs which will severely hinder the effectiveness of this estimator.
Figure 1: Left/Center: Herbie function with failure contour (red) and input distribution (white). Right: Entropy surface (yellow/high, purple/low) for a GP trained on a 100-point LHS (white circles).
Figure 2: Slice of Herbie function along $x_{2}=1.1$. GP was trained on 100-point LHS in 2 dimensions.
For example, consider the two-dimensional “Herbie” function (Lee et al.,, 2011) shown in the left and center panels of Figure 1 (we will revisit the right panel momentarily). We define a failure threshold of $t=1.065$ which demarcates four disjoint failure regions (red contours). We also define $p(\mathbf{x})$ using independent normal distributions centered at the origin (white contours). [^1] Full details of all simulated functions, thresholds, and distributions are provided in the Supplementary Material. We trained a GP surrogate using a random Latin hypercube sample (LHS; McKay et al.,, 2000) of size $n=100$ in two dimensions. Figure 2 shows the GP posterior prediction along the slice $x_{2}=1.1$. The posterior mean (dashed blue) and 95% confidence interval (blue shading) capture the true nonlinear surface (solid black) fairly well. Even so, there are inputs (highlighted by the yellow shading) which the surrogate would misclassify as “passes” instead of “failures.” Thus any estimate reliant solely on this surrogate would underestimate the failure probability.
Deep Gaussian processes.
In some situations, surrogate predictions may be improved through adaptations to the surrogate model itself. Traditional GPs are limited by the commonly assumed stationarity of the covariance kernel; they are not able to accurately model nonstationary surfaces characterized by regime shifts and stark transitions. Deep Gaussian processes (Damianou and Lawrence,, 2013) are a promising more flexible alternative. They leverage functional compositions of GP layers in which latent layers act as warped versions of the original inputs, allowing for greater flexibility. DGPs have consistently outperformed stationary GPs in modeling nonstationary computer experiments (e.g., Rajaram et al.,, 2021; Yazdi,, 2022; Ming et al.,, 2023), particularly when deployed in a Bayesian framework using MCMC sampling to achieve full posterior integration of latent layers (e.g., Sauer et al.,, 2023b, a). See Sauer, (2023) for a thorough review of DGP surrogates. Crucially, our method is agnostic to surrogate choice as long as UQ is provided. We will demonstrate functionality with both traditional GP and Bayesian DGP surrogates in Section 4.
2.2 Contour location
To facilitate failure probability estimation (4), a surrogate must accurately classify passes and failures. Strategic sequential designs targeting the failure contour, i.e., “contour location,” will outperform space-filling counterparts in this regard. In CL, after initializing with a small space-filling design, subsequent training locations are selected iteratively through the optimization of an acquisition criterion. The computer simulation is evaluated at the selected input(s), the surrogate is updated, and the process is repeated until the budget is exhausted or a stopping criterion is met.
Identifying acquisition criteria which effectively target failure contours is a hot topic, particularly in conjunction with GP surrogates, (e.g., Bichon et al.,, 2008; Ranjan et al.,, 2008; Bect et al.,, 2012; Chevalier et al.,, 2014; Marques et al.,, 2018; Cole et al.,, 2023). Here, we highlight one in particular – classification entropy – defined as
$H(\mathbf{x})=-p_{\mathbf{x}}\log(p_{\mathbf{x}})-(1-p_{\mathbf{x}})\log(1-p_{\mathbf{x}}),$ (5)
where $p_{\mathbf{x}}$ represents the probability of failure at input $\mathbf{x}$. With a GP surrogate, this probability boils down to a Gaussian CDF computation:
$p_{\mathbf{x}}=\mathrm{P}\left(\hat{f}_{n}(\mathbf{x})>t\right)=1-\Phi\left(\frac{t-\mu_{n}(\mathbf{x})}{\sigma_{n}(\mathbf{x})}\right).$
Returning to the Herbie example, the entropy surface for the trained GP is shown in the right panel of Figure 1 and along the slice in Figure 2 (green dotted/dashed). Notice the high entropy regions mostly agree with the locations of the true failure contour (we could improve upon this with a CL design instead of an LHS). A reasonable acquisition could fall in any of these local optima.
Naturally, implementations of entropy as an aqcuisition function vary on the finer details. For GP surrogates we will leverage the optimization scheme of Cole et al., (2023), which first deploys a small LHS, identifies the LHS sample with the highest entropy, then runs a numerical optimization of entropy initialized at that point. By avoiding multi-starts, Cole et al., target local optima in the entropy surface. For DGP surrogates, we will follow Booth et al., (2024b). To avoid cumbersome numerical optimizations with the heftier MCMC-based DGP, Booth et al., use triangulation candidates (Gramacy et al.,, 2022), then select the candidate on the Pareto front of entropy and uncertainty. Henceforth, all GPs and DGPs we consider will be trained with these two contour location methods, respectively.
Figure 3: Various designs for estimating the failure probability of the Herbie function with a GP surrogate and a total budget of $N=150$. All methods start with the same CL design of size $n=100$ (black circles), leaving a remaining budget of $B=50$ (blue triangles) for either continued surrogate training or importance sampling. Left: Entire budget is spent on continued contour location, with acquisitions exploring the entire space. Center: Remaining budget is sampled from a bias distribution for importance sampling (black circles omitted since they do not feature in the SIIS estimate). Right: Remaining budget is used for MC samples with highest entropy, solely exploiting near the predicted contour. Comparison of failure probability estimates obtained from these is provided later in Section 4. Although there is no clear distinction in performace between the left/right designs for this simple illustrative example, the continued exploration of the “All CL” design will hinder performance on larger, harder problems (Figure 6).
Continuing the Herbie function example, the left panel of Figure 3 shows the design points chosen from entropy-based contour location with a GP surrogate (Cole et al.,, 2023), indicated by both the black circles and blue triangles (the distinction between the two will become clear when we discuss the other panels). Notice, the chosen points attempt to balance exploration (points spread out over the entire domain) and exploitation (points hovering around the red failure contour). The resulting surrogate may inform an $\hat{\alpha}_{\textrm{SURR}}$ estimate (4); we will include this as a competitor in our later exercises.
2.3 Importance sampling
Rather than side-stepping the prohibitive computation of Eq. (1) with surrogate evaluations, importance sampling (Tokdar and Kass,, 2010) avoids expensive evaluations by sampling from a different distribution altogether. Let $q(\mathbf{x})$ denote a “bias” density which shares domain with $p(\mathbf{x})$. The failure probability $\alpha$ may be equivalently represented as
[…]
Monte Carlo approximation of this integral provides the estimator
[…]
The “bias weights” $w_{i}$ account for the difference between the input density and the bias density. This estimator is theoretically unbiased and requires a smaller budget of simulator evaluations ($B\ll M$) if the bias distribution appropriately covers and targets the failure region (Srinivasan,, 2002). But estimation of an effective bias distribution is tricky. The ideal bias distribution is $q(\mathbf{x})^{*}=\frac{1}{\alpha}\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}p(\mathbf{x})$, but neither $\alpha$ nor the indicator function are known a priori. For the sake of feasibility, $q(\mathbf{x})$ is traditionally constrained to a known distributional family, such as a Gaussian mixture model (GMM; Reynolds et al.,, 2009). Bias distribution training is then relegated to learning distributional parameters (in a GMM, these include the number of mixtures, mixture weights, and component means and variances), often accomplished through cross validation or information criteria maximization.
In “surrogate-informed importance sampling,” Peherstorfer et al., (2016) leverage a surrogate to identify predicted failures from a large Monte Carlo sample of the input distribution, then train a GMM bias distribution to these predicted failures. If the surrogate effectively identifies the failure region, then these samples should mimic the ideal $q(\mathbf{x})^{*}$. With a trained bias distribution in-hand, samples may be drawn from $q(\mathbf{x})$, evaluated through the expensive simulator, and deployed in Eq. (6) to provide an estimate of the failure probability.
To illustrate, consider our running Herbie function example. Using the GP surrogate trained on the 100-point contour locating design from Figure 3 (black circles) and a large MC sample from the input distribution, we trained a GMM bias distribution to the MC samples which the surrogate predicted to be failures. The blue triangles in the center panel show 50 points sampled from this bias distribution. These samples may be evaluated through the true model to estimate $\hat{\alpha}_{\textrm{IS}}$ (6), thereby exhausting our entire budget of 150 evaluations. While the blue triangles target the failure region, they do not provide enough coverage to ensure a precise estimate. In our later experiments (Section 4), SIIS performance is highly variable.
While SIIS has been widely embraced for its unbiasedness, it has several key weaknesses. First, the bias distribution, which features heavily in the final estimate of Eq. (6), is a completely auxilary quantity. Hypothetically, two different bias distributions could produce the same samples $\left\{\mathbf{x}_{i}\right\}_{i=1}^{B}$, thus producing the same $\left\{f(\mathbf{x}_{i})\right\}_{i=1}^{B}$. But in the calculation of Eq. (6), these two bias distributions would contribute different weights, resulting in different final answers even though the training data from the computer simulator was exactly the same. O’Hagan, (1987) first pointed out this conundrum as a violation of the “Likelihood principle.” A second weakness is the restriction of $q(\mathbf{x})$ to distributional families like GMMs. Failure regions may be complex and may not fit nicely under these target distributions. Consequently, SIIS typically requires large $B$ in order to adequately cover the failure region. For context, Peherstorfer et al., (2016) entertained $B\in\left\{100,1000,10000,100000\right\}$ and found performance continued to improve as $B$ increased. Furthermore, there is no clear avenue for determining the budget allocation between surrogate training and importance sampling estimation. [Throughout, we will borrow from our proposed method to inform budget allocation for the SIIS competitor, which is how we chose to separate the budget of Figure 3 into 100 for training and 50 for estimation.] Finally, the SIIS method does not leverage the full potential of the surrogate model. Many expensive simulator evaluations go into training the surrogate and, in the end, the surrogate is only used to identify predicted failures, with no mention of its UQ. [Notice, as portrayed in the center panel of Figure 3, the black circles are abandoned after the bias distribution is trained.]
3 Two-stage design for failure probability estimation
All in all, we are unsatisfied with the two state-of-the-art surrogate-based methods for estimating $\alpha$ which we just described: surrogate MC and SIIS. While the drawbacks to SIIS are apparent, the limitations of surrogate MC deserve more attention. To investigate further, we conducted the same CL sequential design of the Herbie function (Figure 3), this time storing the surrogate MC estimate $\hat{\alpha}_{\textrm{SURR}}$ (4), which we now denote as simply $\hat{\alpha}$, after every 10th acquisition. Progress is shown in the left panel of Figure 4. While there is significant learning occurring with early acquisitions, it quickly plateaus. By the time the surrogate has been trained on 100 points, it knows just as much about $\alpha$ as it does when it has been trained on the full budget of $N=150$ points. In this case, there is bound to be a more effective way to use the remaining simulation budget. We propose a solution: first developing a stopping criterion to decide when to halt contour location, then proposing a hybrid Monte Carlo estimator that greedily leverages the remaining simulation budget precisely for failure probability estimation.
Figure 4: Left: Surrogate MC failure probability estimate for the Herbie function, updated after every 10th contour locating acquisition. True $\alpha$ in dashed red. Right: Difference in successive estimates with ±$\pm\hat{\sigma}_{\alpha}$ shown in dashed green. Blue square marks the second update within ±$\pm\hat{\sigma}_{\alpha}$.
3.1 A stopping criterion for contour location
Let $n$ denote the budget used on contour location and $B$ denote the remaining budget such that the total budget is $N=n+B$. We seek an objective, automatable decision rule for selecting $n$. The right panel of Figure 4 shows the relative difference between each successive $\hat{\alpha}$ estimate from the left panel. As expected, the relative differences converge to zero as the surrogate hones in on the true contour, but they never settle at exactly zero. How small of an update then is small enough to indicate learning has plateaued?
The significance of an update to $\hat{\alpha}$ is only interpretable within the context of $\alpha$, the true failure probability, and $M$, the size of the MC sample used. Small differences might be inconsequential for large $\alpha$ with small $M$, but could be significant for small $\alpha$ with large $M$. To this end, we propose the use of the standard error $\sigma_{\alpha}$ to quantify the relative magnitude of each update. Notice in the estimator of Eq. (2), the indicator $\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}$ is a Bernoulli random variable with expected value $\alpha$ and variance $\alpha(1-\alpha)$. Thus, the MC estimator has variance inversely proportional to $M$:
$\mathbb{V}[\hat{\alpha}_{\textrm{MC}}]=\frac{1}{M^{2}}\sum_{i=1}^{M}\mathbb{V}\left[\mathbbm{1}_{\left\{f(\mathbf{x})>t\right\}}\right]=\frac{\alpha(1-\alpha)}{M}.$
Let $\sigma_{\alpha}=\sqrt{\frac{\alpha(1-\alpha)}{M}}$ denote the standard error of this estimator. This standard error intrinsically accounts for the magnitude of $\alpha$ and $M$. Since the true $\alpha$ is unknown, we estimate $\hat{\sigma}_{\alpha}=\sqrt{\frac{\hat{\alpha}(1-\hat{\alpha})}{M}}$ using the current surrogate MC estimate.
The green dashed lines in the right panel of Figure 4 show ±$\pm\hat{\sigma}_{\alpha}$, updated along with $\hat{\alpha}$. At the start, updates to $\hat{\alpha}$ are far beyond one standard error. But after a while, the updates sit comfortably within these bounds. Let $X_{M}$ denote the row-combined matrix of $M$-many samples from $p(\mathbf{x})$. We posit that, for fixed $X_{M}$, any update to $\hat{\alpha}$ within ±$\pm\hat{\sigma}$ is “small enough.” Accordingly, we propose halting CL once we have observed two successive updates within ±$\pm\hat{\sigma}$. The blue square in Figure 4 marks this point, stopping CL for this example at $n=100$, effectively identifying a point where surrogate learning has saturated and leaving a remaining budget of $B=50$ (this informed the budget allocation of Figure 3).
Note, this stopping criterion requires the same $X_{M}$ be used in every $\hat{\alpha}$ calculation. If a new $X_{M}$ is generated for each estimate, there will be additional variability among the $\hat{\alpha}$’s, and the proposed ±$\pm\hat{\sigma}_{\alpha}$ threshold would need to be increased. Removing this extraneous source of variability allows successive estimates to pinpoint differences stemming solely from the surrogate.
Our budgeting scheme is summarized as follows. Initialize a surrogate, begin a contour locating sequential design, and collect one sample $X_{M}$. At the start, it is likely that there will be no failures observed and that $\hat{\alpha}$ will be equal to zero (as seen in Figure 4). To safeguard against halting surrogate training prematurely, we build in a “common sense” check by setting a minimum number of failures that must be observed before the contour location could be halted. If the contour location design is succeeding, it will be placing some points in the failure region. We recommend a default of 10 failures, but suggest increasing this for more complex, higher dimensional problems. Once the design has surpassed these minimum requirements, estimate $\hat{\alpha}$ after every 10th acquisition. Calculate $\hat{\sigma}_{\alpha}$ and check if $|\hat{\alpha}^{\textrm{(current)}}-\hat{\alpha}^{\textrm{(previous)}}|<\hat{\sigma}_{\alpha}$. When this condition has been satisfied twice in a row, halt the sequential design and proceed to the next stage (the topic of Section 3.2).
Choosing M.
In practice, it is important to select a MC sample size that is large enough without being too cumbersome. Larger $M$ will produce more effective sum approximations of the true integrals, but we need surrogate predictions (e.g., Eq. 3) for every sample in $X_{M}$. Although this computation is significantly cheaper than evaluation of the expensive simulator, it may still be significant for extremely large $M$, particularly with a heftier surrogate like a Bayesian DGP. If we knew $\alpha$, we could determine $M$ by thresholding the standard error $\sigma_{\alpha}$. For example, to achieve a standard error less than 10% of $\alpha$ requires:
$\sigma_{\alpha}<\frac{\alpha}{10}\quad\longrightarrow\quad\frac{100(1-\alpha)}{\alpha}<M.$
To achieve this precision, $\alpha=10^{-4}$ requires $M$ nearly 1 million, $\alpha=10^{-5}$ requires $M$ nearly 10 million, and $\alpha=10^{-6}$ requires $M$ nearly 100 million. In practice, with unknown $\alpha$, we recommend either choosing $M$ conservatively given expert insight into the anticipated failure probability or performing a pilot study where $M$ is incremented say by one million at a time. Gather surrogate predictions along this incremental process, and use $\hat{\alpha}$ to inform $\hat{\sigma}_{\alpha}$.
3.2 Hybrid Monte Carlo for failure probability estimation
Once $n$ is determined, how should we spend the remaining budget of simulator evaluations ($B=N-n$), in conjunction with our trained surrogate, to most effectively estimate $\alpha$? We propose an upgrade to the surrogate MC estimator of Eq. (4). Let $\mathcal{U}\subset\mathcal{X}$ represent some subset of the input domain (boundary inclusive). Then, Eq. (1) may be partitioned as
[…]
inspiring an equivalent formulation of the surrogate MC estimator,
[…]
Now for $\mathbf{x}_{i}\in\mathcal{U}$, upgrade surrogate prediction $\mu_{n}(\mathbf{x}_{i})$ to expensive evaluation of the true $f(\mathbf{x}_{i})$, resulting in the following “hybrid MC” estimator:
[…]
The hybrid nature of this estimator is in the use of both true simulator evaluations (left term) and surrogate evaluations (right term). The accuracy of this estimator is contingent upon the classification accuracy of the surrogate and the choice of $\mathcal{U}$. As long as we restrict the size of $\mathcal{U}$ so only $B$-many samples from $X_{M}$ fall within it, i.e., $\sum_{i=1}^{M}\mathbbm{1}_{\left\{\mathbf{x}_{i}\in\mathcal{U}\right\}}=B$, then we will meet our budgetary constraints.
Rather than thinking of $\mathcal{U}$ as a continuum, we find it easier to conceptualize in the context of the discrete set $X_{M}$. In a perfect world, $\mathcal{U}$ would capture all the samples which the surrogate would misclassify. We are not privy to this information, and we are under a limited budget, but we can leverage the surrogate’s UQ to identify the $B$-many samples at which it is most uncertain. Recall, classification entropy (5) quantifies the degree of uncertainty in the pass/fail prediction of a surrogate. Consequently, our proposed allocation of samples to $\mathcal{U}$ is strikingly simple – place the $B$-many samples from $X_{M}$ with the highest entropy into $\mathcal{U}$. Denote these as $X_{M}^{B}$. Revisiting Figure 2, notice the high entropy regions (green dotted/dashed) align well with the misclassification regions (yellow shading). The entropy criterion is effectively identifying regions of high classification uncertainty where we could most utilize expensive true simulator evaluations to correct inaccurate surrogate predictions.
Ultimately, after evaluating $f(\mathbf{x})$ for $\mathbf{x}\in X_{M}^{B}$, we update the surrogate with these additional observations, thereby updating $\mu_{n}(\mathbf{x})$, before returning $\hat{\alpha}_{\textrm{HYB}}$ (7). Since the $X_{M}^{B}$ points are likely near a failure contour, they could significantly improve the accuracy of the surrogate predictions for other samples near the contour. For reference, consider the right panel of Figure 3. The black circles showing the contour locating design for the GP surrogate are unchanged, but here the additional budget of 50 evaluations is spent on the MC samples from $p(\mathbf{x})$ with the highest classification entropy (blue triangles). Compared to the left panel which continues to explore the entire domain, these highest entropy samples exclusively exploit the predicted failure contour. Compared to the bias samples portrayed in the middle panel, where the original surrogate design is abandoned, our two-stage design is able to leverage all simulator evaluations in its final estimate.
3.3 Summary and related work
We have presented a comprehensive strategy for estimating Eq. (1) from a budget of size $N$. It is summarized as follows. Start with a very large Monte Carlo sample from $p(\mathbf{x})$, denoted $X_{M}$. Train a surrogate model through a contour locating sequential design, monitoring the progress of the $\hat{\alpha}_{\textrm{SURR}}$ estimate for given $X_{M}$, and halting contour location once two consecutive updates are within ±$\pm\hat{\sigma}_{\alpha}$. Denote the design points chosen in this stage as $X_{n}$. Then, use the trained surrogate to identify the highest entropy samples, $X_{M}^{B}$, evaluate the simulator at these inputs, and update the surrogate with these new observations. Finally, return $\hat{\alpha}_{\textrm{HYB}}$ following Eq. (7).
Given the deterministic nature of $f(\mathbf{x})$, effective surrogates should interpolate observed points. In this light, the hybrid MC estimator (7) is akin to a surrogate MC estimator (4) when the surrogate has been trained on all $\mathbf{x}_{i}\in\mathcal{U}$. This connection permits us to frame our contribution as a two-stage design. In the first stage, $X_{n}$ is chosen over $\mathcal{X}$ based on contour location. In the second stage, $X_{M}^{B}$ is chosen over $\mathbbm{P}_{\mathbf{x}}$ based on maximum entropy. The final design of size $N$ is then $X_{N}=[X_{n}^{\top},{X_{M}^{B}}^{\top}]^{\top}$.
There are some noteworthy distinctions between these two design stages. To illustrate, refer back to the black circles ($X_{n}$) and blue triangles ($X_{M}^{B}$) in the right panel of Figure 3. In the first stage, the $n$ points selected for contour location may be acquired anywhere in the domain $\mathcal{X}$, seeking to balance both exploration of the response surface and exploitation of the suspected contour. In fact, the contour locating sequential design strategies we consider in Section 4 (Cole et al.,, 2023; Booth et al.,, 2024b) intentionally circumvent acquisitions based on maximal entropy in order to avoid clustered acquisitions and to promote more exploration of the surface. Alternatively, the $B$ points chosen in the second stage must originate from $p(\mathbf{x})$ and are solely exploitative. By nature, they will be clustered around the predicted contour. We contend that having a mix of data from these two realms is superior to either extreme.
Some of the works we referenced in Section 1 have investigated stopping criteria for surrogate training in reliability settings. The most similar to our proposed criterion is that of Zhu and Du, (2016) who recommend a convergence criterion akin to $\frac{\hat{\sigma}_{\alpha}}{\hat{\alpha}}\leq\epsilon$, where $\epsilon$ is a user specified threshold. While similar to our stopping criterion, the choice of $\epsilon$ here is make-or-break, and little guidance is provided. If $\epsilon$ is too high, contour location will be halted too soon; if too low, contour location will continue past the point of utility. Our proposed stopping criterion avoids this choice and may be easily automated without problem-specific tuning. In our synthetic experiments, we find that it consistently results in appropriate termination with budget to spare.
Hybrid MC estimators (similar to Eq. 7) have been entertained in other settings. Li and Xiu, (2010) use polynomial chaos models for a hybrid MC estimator with samples delegated to $\mathcal{U}$ based on predicted proximity to the contour, i.e., $|\mu_{n}(\mathbf{x}_{i})-t|$. This work, however, does not consider surrogate design and does not leverage surrogate UQ. Fuhrländer and Schöps, (2020) use hybrid MC, but they only consider training the surrogate on samples from $p(\mathbf{x})$. Echard et al., (2011) similarly train only on a subset of $X_{M}$. We attempted to recreate the approach of Fuhrländer and Schöps, (2020) in our setting: (i) sample large $X_{M}$, (ii) train an initial surrogate on a random selection of these samples, (iii) use the surrogate with a metric such as entropy to select additional batches from $X_{M}$ for evaluation through the simulator, (iv) update the surrogate, and (v) repeat steps (iii)-(iv) until the budget is exhausted. Yet we were unable to obtain even reasonable results because initial surrogate fits were very poor, and surrogate-informed acquisitions were unable to recover. Random selections from $X_{M}$ were overwhelmingly in high probability regions, not low probability failure regions. Hristov et al., (2019) noticed this conundrum and stated it well: “using a surrogate that is built on frequent events to predict and analyze rare events” can be futile.
To visualize this conundrum in a simple setting, we revisit the Herbie example of Figure 1. With the input distribution centered at the origin, the probability of generating any samples outside $[-1,1]^{2}$ is very low. But a GP surrogate would benefit from some “anchor points” in these outer regions (Gramacy and Apley,, 2015). This distinction will be less prominent if $p(\mathbf{x})$ is uniformly distributed over $\mathcal{X}$, but the trade-off between exploring in the training phase and exploiting in the estimation phase will still be impactful. Even with uniform $p(\mathbf{x})$, training a surrogate only on maximal entropy points is suboptimal (Cole et al.,, 2023; Booth et al.,, 2024b).
4 Synthetic experiments
In this section we benchmark our method against state-of-the-art competitors on a variety of synthetic experiments. We consider four functions: the 2d Herbie function (Lee et al.,, 2011) with a total budget of 150, the 3d Ishigami function (Ishigami and Homma,, 1990) with a total budget of 300, the 6d Hartmann function (Picheny et al.,, 2013) with a total budget of 600, and the 4d Plateau function (Booth et al.,, 2024b) with a total budget of 200. The first three of these are plausibly stationary, so we use a standard GP surrogate, fit with the Scikit-learn package in python (Pedregosa et al.,, 2011), and the entropy contour location scheme of Cole et al., (2023). The Plateau function is known to be nonstationary; it is characterized by flat regions with a steep sloping drop between them. To address the nonstationarity, we use a Bayesian DGP surrogate (Sauer et al.,, 2023b), fit with the deepgp package in R (Booth,, 2023), with the contour location scheme of Booth et al., (2024b). With the right surrogate, the failure contour of the Plateau function can be well estimated from a rather small amount of data.
Function True $\alpha$ $d$ $n_{0}$ $N$ $M$ Surrogate CL Scheme
Herbie $7.533\times 10^{-5}$ 2 20 150 $3.5\times 10^{7}$ GP Cole et al., (2023)
Ishigami $1.904\times 10^{-4}$ 3 50 300 $1.5\times 10^{7}$ GP Cole et al., (2023)
Hartmann $1.001\times 10^{-5}$ 6 100 600 $1.0\times 10^{8}$ GP Cole et al., (2023)
Plateau $4.308\times 10^{-4}$ 4 30 200 $3.5\times 10^{6}$ DGP Booth et al., (2024b)
Table 1: Simulation settings. True $\alpha$ values are estimated from 10 billion samples.
Table 1 summarizes the simulation settings. Functions are observed without noise to mimic a deterministic computer simulation. We define failure thresholds and input distributions for each function, resulting in the failure probabilities reported in Table 1. All input distributions are variations/combinations of independent uniform and truncated normal distributions; further details of the functions, thresholds, and distributions are reserved for Supplementary Material. We choose $M$ for each $\alpha$ to keep $\sigma_{\alpha}$ conservatively under 5%.
Figure 5: Distribution of the chosen $n$ across 30 repetitions.
First, we deploy our stopping criterion (Section 3.1) for each of the four scenarios, starting with an initial LHS sample of size $n_{0}$ (Table 1) and repeating for 30 re-randomized Monte Carlo repetitions. To avoid premature termination, we require a minimum of 10 observed failures and $n\geq 2n_{0}$ (atleast as many contour locating acquisitions as original space-filling observations). Figure 5 shows the distribution of chosen $n$ values. Notice, all trials terminated before the allotted $N$, reserving some evaluations for additional estimation, but there was considerable variability across trials. We view this variability as a strength – some random initializations need more data to effectively locate the contour while some get lucky from the start. Allowing the surrogate to dictate its own budget mitigates some of the risks of having an unlucky starting design.
For each trained surrogate with chosen $n$ (which varies across repetitions), we collect a large Monte Carlo sample ($X_{M}$, randomized across repetitions), then deploy the following estimation methods:
  • All CL: Ignoring the chosen $n$, continue contour location until the entire budget is exhausted and return the surrogate MC estimate (4).
  • Hybrid Entropy: Our two-stage approach using the hybrid MC estimate (7) where $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ with highest entropy.
  • Hybrid Proximity: Same as “Hybrid Entropy,” but $\mathcal{U}$ is chosen to contain the $B$ points from $X_{M}$ which are predicted to be closest to the failure threshold (i.e., the samples with the smallest $|\mu_{n}(\mathbf{x}_{i})-t|$). This is inspired by the work of Li and Xiu, (2010).
  • SIIS: Surrogate-informed importance sampling (6) with a GMM bias distribution (Peherstorfer et al.,, 2016). The bias distribution is trained on the predicted failures, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})>t\right\}$. For our GP surrogates implemented in python, we use the scikit-learn package (Pedregosa et al.,, 2011) to train GMM bias distributions. For our DGP surrogates implemented in R, we use the mclust package (Scrucca et al.,, 2016).
  • SIIS UCB: Same as “SIIS,” but the bias distribution is trained on the samples whose upper 95% confidence bound exceeds the threshold, i.e., $\left\{\mathbf{x}_{i}\in X_{M}\;|\;\mu_{n}(\mathbf{x}_{i})+1.645\sigma_{n}(\mathbf{x}_{i})>t\right\}$.
Each method uses the same $X_{M}$ and the same surrogate trained up to point $n$. The distinction is in what is done with the remaining $B$ samples. As implemented, the latter three are to some degree an internal comparison since they use our proposed stopping criterion to determine the budget allocation between surrogate training and estimation (Figure 5). Without borrowing from our proposed approach, there would be no guidance on the allocation of $n$ and $B$ for these methods. The “Hybrid Proximity” approach is truly an internal comparison as it borrows everything from our proposed two-stage design aside from the specific allocation of samples to $\mathcal{U}$. It is included to show the impact of the entropy selection criterion, which uses surrogate UQ, over a selection solely based on the predicted mean. Reproducible code for all experiments, including both the python and R implementations, is available in our public git repository. [^2] https://bitbucket.org/boothlab/failprob/
Figure 6: Estimated failure probabilities shown over true failure probability (red dashed) ±$\pm 2$ standard errors (grey shading). Boxplots show 30 repetitions.
Results are shown in Figure 6. Red dashed lines mark the true failure probabilities, and grey shaded regions highlight $\{\alpha\pm 2\sigma_{\alpha}\}$ where $\sigma_{\alpha}$ is calculated from the true $\alpha$ and chosen $M$. This region provides important context – it captures the intrinsic variability that would accompany any MC estimator. In other words, even if we used the true simulator for all $\{\mathbf{x}_{i}\}_{i=1}^{M}$ samples (2), we would obtain a boxplot roughly spanning the grey interval. We thus consider any results within this region to be effective estimates. We could tighten the interval (and the boxplots of the MC methods) by increasing $M$ without requiring any additional simulator evaluations, another advantage MC holds over IS.
Across the board, both importance sampling based methods (“SIIS” and “SIIS UCB”) suffered from high variability in their failure probability estimates. Although their median performance was generally on target (indicative of their unbiasedness), these methods consistently offered the poorest performance. The more conservative “SIIS UCB” had no apparant advantage over the traditional “SIIS.” We believe importance sampling is simply ill-suited for our low data settings. In all scenarios aside from the simple two-dimensional Herbie function, our two-stage design (“Two-stage Entropy”) outperformed the exhaustive contour location approach (“All CL”). Particularly as dimension grows and the response surface becomes more complex, there is a clear advantage to reserving some simulator evaluations for the second stage. Most of the time there was agreement between the two criteria for allocating samples to $\mathcal{U}$ (“Entropy” and “Proximity”), but when there was disagreement the entropy criterion offered higher accuracy, with uniformly superior worst-case performance. As suspected, the entropy criterion which incorporates uncertainty is more effective at identifying samples which the surrogate would misclassify.
5 RAE-2822 airfoil computer experiment
The RAE-2822 airfoil computer experiment simulates the flow of air around an aircraft wing. The simulation suite is solved via SU2, and the software is publicly available (Economon et al.,, 2016). We consider seven inputs: four shape parameters ($S_{i}$ for $i=1,\dots,4$) and 3 environmental parameters, namely angle of attack, Reynolds number, and Mach number. We are interested in the lift over drag ($L/D$) ratio. $L/D$ values that are too low represent inefficient aircraft performance. We define a failure threshold at $t=3$ such that $L/D<3$ indicates failure. [Note the change from $f(\mathbf{x})>t$ to $f(\mathbf{x})<t$ does not affect the implementation of our method.] This computer experiment requires nearly 30 minutes of compute time with serial execution, but reductions are possible with parallel execution.
Our objective is to estimate the probability of failure given the following input distribution:
$S_{i}$ $\sim\mathrm{Unif}(-1\times 10^{-4},1\times 10^{-4})\;\;\textrm{for}\;\;i=1,\dots,4$
Angle of Attack $\sim\mathcal{N}\left(\mu=5,\sigma=1\right)\;\;\textrm{truncated to}\;\;[0,10]$
Reynolds number $\sim\mathcal{N}\left(\mu=1\times 10^{7},\sigma=1\times 10^{6}\right)\;\;\textrm{truncated to}\;\;[5\times 10^{6},1.5\times 10^{7}]$
Mach number $\sim\mathcal{N}\left(\mu=0.8,\sigma=0.02\right)\;\;\textrm{truncated to}\;\;[0.7,0.9]$
We fix our total budget at $N=500$ simulator evaluations and use $M=2.5\times 10^{6}$ for the size of $X_{M}$. Booth et al., (2024b) found that DGP surrogates outperformed GP surrogates for contour location of this experiment. We follow the settings therein, starting with an LHS sample of size $n_{0}=100$ and fitting a Bayesian DGP surrogate with the deepgp package. We proceed with contour location following Booth et al., (2024b), stopping after every 10th acquisition to estimate $\hat{\alpha}$ and check whether it is with in the ±$\pm\hat{\sigma}_{\alpha}$ bounds. Once we halt contour location (Section 3.1), we use entropy to inform our hybrid MC estimator (Section 3.2). Given the larger input dimension, we require at least 40 observed failures before we consider halting surrogate training. Since the true failure probability is unknown, we opt to repeat this entire exercise twice, with a newly randomized starting design. We consider the agreement between the two randomized repetitions as an indication that our procedure is well-calibrated.
Figure 7: Progress in $\hat{\alpha}$ across two contour locating designs for the airfoil computer experiment. Proposed two-stage failure probability estimate shown in dashed red. Grey shading indicates $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$ for these final estimates. Left/right panels show two repetitions with newly randomized starting designs.
Figure 7 reports the progress in $\hat{\alpha}$ over the course of each design. Our stopping criterion halted the sequential designs at $N=180$ and $N=160$, leaving $B=320$ and $B=340$ for the second design stage, respectively. Our two hybrid Monte Carlo estimates were $\hat{\alpha}_{\textrm{HYB}}=\{0.0003368,0.0003432\}$. These values are indicated by the red dashed lines in Figure 7. The grey shading shows $\hat{\alpha}\pm 2\hat{\sigma}_{\alpha}$, calculated from these final estimates. Although the two estimates vary slightly, they are both within the bounds of reasonable variability for our chosen $M$. Additionally, notice how the final hybrid MC estimate of the second exercise (right panel) brought the failure probability lower than what the surrogate had originally estimated, closer to the estimate of our first attempt (left panel). We take this as an indication that the hybrid estimator is effectively adjusting for surrogate inaccuracies.
6 Discussion
We presented a unique two-stage surrogate design for estimating failure probabilities of expensive computer experiments. Starting with a contour locating sequential design, we halt contour location once surrogate training has plateaued, as indicated by practical convergence of a surrogate Monte Carlo estimator. Then we use the remaining budget of simulator evaluations on the Monte Carlo samples with highest entropy, to ultimately return a hybrid MC estimate of $\alpha$. Our design strategically leverages the uncertainty quantification provided by the surrogate to reserve expensive simulator evaluations for samples with high classification uncertainty. While we limited our experiments to GP and DGP surrogates, our proposed scheme is applicable to any surrogate which provides posterior predictive distributions (which are required for entropy calculations). Extensions to other surrogates, such a treed Gaussian processes (Gramacy and Lee,, 2008) or Bayesian additive regression trees (Chipman et al.,, 2010), is straightforward.
Our two-stage design is computationally efficient in the number of evaluations of the expensive simulator, but it requires many predictive evaluations from the surrogate. Posterior predictions must be computed at all $M$ samples multiple times. While surrogate predictions are very cheap, collecting billions of them is not necessarily trivial. Our experiments were limited to failure probabilities above $1\times 10^{-5}$ for this very reason. We could tackle smaller probabilities, but it would require larger $M$, which would eventually become impractical. Although we are focused on expensive simulators whose computational costs overshadow the cost of these surrogate predictions, our approach may simply not be suited to extremely rare failure probabilities.
In all of our examples, our stopping criterion terminated contour location before the complete budget ($N$) was exhausted. If a simulation budget is severely restricted, the budget could be exhausted before our proposed stopping criterion is met. While we prefer to have additional evaluations reserved for the estimation stage, we contend that using the entire budget for contour location to provide a surrogate MC estimate is the best option in this scenario. As we’ve shown, importance sampling based methods perform poorly when data is limited.
We have exclusively focused on deterministic computer experiments. In fact, the noise-free nature of the simulator is what enabled us to frame our contribution as a two-stage design. Once $f(\mathbf{x}_{i})$ has been observed and incorporated into the surrogate, surrogate predictions should interpolate these observations, i.e., $\mu_{n}(\mathbf{x}_{i})=f(\mathbf{x}_{i})$. Thus our hybrid MC estimator (7) is really a surrogate MC estimator (4), with a strategically trained surrogate. Different design choices could be warranted for stochastic simulators (Baker et al.,, 2022). In particular, the hybrid estimator of Eq. (7) should not make use of noisy $f(\mathbf{x})$. Replicates in the design may be crucial (Binois et al.,, 2019).
When the input distribution is not uniform over $\mathcal{X}$, it could be beneficial to leverage information about $p(\mathbf{x})$ in the sequential training of the surrogate (Abdelmalek-Lomenech et al.,, 2024). There is arguably no need to train a surrogate in a region where inputs will never occur. Incorporating this information in the contour location stage could improve failure probability estimation downstream, but we have seen positive impacts from having “anchor points” in regions of very low density (such as the black open circles around the edges of Figure 3). We suspect there is a delicate balance that would work well here, but this avenue is currently underexplored.
Estimation of failure probabilities is often an intermediate step within a larger objective, such as a stochastically constrained optimization. For example, in the context of our airfoil experiment, we may wish to find the optimal wing configuration given that the probability of failure remains below a specified threshold. The ability to effectively estimate failure probabilites from limited data, which we have tackled here, is integral to the success of this larger enterprise.
SUPPLEMENTARY MATERIAL
Here we provide the details of the functions, thresholds, and input distributions used in the benchmark exercises of Section 4. In each case, failures are defined as $f(\mathbf{x})>t$.
Herbie.
The 2d Herbie function (Lee et al.,, 2011) is defined in over $[-2,2]^{2}$ as
$f(\mathbf{x})=\prod_{i=1}^{2}\mathrm{exp}\left(-(x_{i}-1)^{2}\right)+\mathrm{exp}\left(-0.8*(x_{i}+1)^{2}\right)-0.05*\sin(8*(x_{i}+1)).$
We set the failure threshold at $t=1.065$ and define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0,0.36)$ for $i=1,2$, truncated to $[-2,2]$.
Ishigami.
The 3d Ishigami function (Ishigami and Homma,, 1990) is defined over $[-\pi,\pi]^{3}$ as
$f(\mathbf{x})=\sin(x_{1})+5*\sin(x_{2})^{2}+0.1*x_{3}^{4}*\sin(x_{1}).$
We set the failure threshold at $t=10.244$ and define the input distribution as
$x_{1}$ $\sim\mathcal{N}(-1,1)\;\;\textrm{truncated to}\;\;[-\pi,\pi]$
$x_{2}$ $\sim\mathcal{N}(1.5,1.5)\;\;\textrm{truncated to}\;\;[-\pi,\pi]$
$x_{3}$ $\sim\mathrm{Uniform}(-\pi,\pi).$
Hartmann.
The 6d Hartmann function (Picheny et al.,, 2013) is defined over $[0,1]^{6}$ as
$f(\mathbf{x})=-\sum_{i=1}^{4}\alpha_{i}\mathrm{exp}\left(-\sum_{j=1}^{6}A_{ij}(x_{j}-P_{ij})^{2}\right)\;\;\textrm{where}$
$\alpha^{\top}=\begin{bmatrix}1\\ 1.2\\ 3\\ 3.2\end{bmatrix}\quad\quad A^{\top}=\begin{bmatrix}10&.05&3&17\\ 3&10&3.5&8\\ 17&17&1.7&.05\\ 3.5&.1&10&10\\ 7&8&17&.1\\ 8&14&8&14\end{bmatrix}\quad\textrm{and}\quad P^{\top}=\begin{bmatrix}.1312&.2329&.2348&.4047\\ .1696&.4135&.1451&.8828\\ .5569&.8307&.3522&.8732\\ .0124&.3736&.2883&.5743\\ .8283&.1004&.3047&.1091\\ .5886&.9991&.6650&.0381\end{bmatrix}.$
We set the failure threshold at $t=2.46$ and define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.5,0.1)$ for $i=1,\dots,6$, truncated to $[0,1]$.
Plateau.
The Plateau function (Booth et al.,, 2024b) is defined in arbitrary dimension over $[0,1]^{d}$ as
$f(\mathbf{x})=2*\Phi\left[\sqrt{2}\left(-4-3\sum_{i=1}^{3}(4x_{i}-2)\right)\right]-1$
where $\Phi$ is the standard Gaussian CDF. We use $d=4$ and set the failure threshold at $t=0$. We define the input distribution as $x_{i}\stackrel{{\scriptstyle\mathrm{ind}}}{{\sim}}\mathcal{N}(0.6,0.11)$ for $i=1,\dots,4$, truncated to $[0,1]$.